# escapeinside=||,style=xcode

### v1

@@ Abstract
Keywords:
Question Answering SPARQL Knowledge Graphs.
@@ 1 Introduction
{minted}
sparql SELECT ?result WHERE ?result p:P39 ?m . ?m ps:P39 wd:Q17989863 ; pq:P580 ?start . OPTIONAL ?m pq:P582 ?end . FILTER (?start <= "2011-12-31"8sd:date & (!BOUND( ?end ) || ?end >= "2011-12-31"8sd:date ))
@@ 3.1 GRASP
With GRASP, the generation of SPARQL queries relies on a LLM that interacts with a knowledge graph autonomously by interleaving function calls to search or query the knowledge graph with reasoning about previous and future steps. Note, that this is in style similar to ReAct prompting [49], though we also allow multiple successive reasoning steps or function calls. Starting point of a generation is the GRASP instruction, which contains the name and endpoint of all available knowledge graphs, details about the step-by-step approach, and additional rules to be followed. Note, that this instruction is generic and does not contain any graph-specific information other than their names and endpoints. Optionally, few-shot examples (see Section 3.5) and a feedback mechanism for refining the output (see Section 3.6) can be added. See Fig. 1 for an overview over GRASP’s approach by means of an example for DBLP.
@@ 3.3 Function Sets
We evaluate different variants of our approach, where each variant has a particular subset of the functions from Section 3.2 available. We call such a subset a function set. We start with the basic function set $B=\{{\color[rgb]{0,0.00390625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.00390625,0.28515625}\texttt{ANS}},{\color[rgb]{0,0.00390625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.00390625,0.28515625}\texttt{CAN}},{\color[rgb]{0,0.00390625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.00390625,0.28515625}\texttt{EXE}}\}$, which allows the model to interact with the knowledge graph solely by executing SPARQL queries. For the other functions sets, we build upon $B$ and add LST to enable the model to retrieve RDF triples, and one or more search functions to enable the model to search the knowledge graph. We use SEN and SPR for the function set $S$, SAC for the function set $S_{A}$ and SCN for the function set $S_{C}$. In theory, the model can perform the most sophisticated searches with SAC, followed by SCN, and then the search functions in $S$. The first two can dynamically constrain searches to subgraphs, with SAC providing even more flexibility than SCN because it allows for full-blown SPARQL queries as constraining element rather than a single RDF triple. However, during initial experiments with these functions, we noticed that the model often struggles to use them properly, or only uses them in their most basic form for unconstrained search. Therefore, we introduce the $S_{E}$ function set to strike a balance between ease of use and functionality. This function set builds upon the function set $S$ but contains additional specialized functions for the two most common types of searches a human expert performs when writing SPARQL queries: SPE to look for properties of a given entity, and SOP to look for entities and literals reachable by a given property. For few-shot evaluations we also add one of FSE or FEX to each function set. See Section 3.5 for more details. See Table 1 for an overview over all function sets.
@@ 3.4 Search Indices
The search functions are all based on pre-computed index data structures, which we describe in the following. See Table 3 for details about all used indices.
@@ 3.4.1 Data
The data for the index data structures comes from two SPARQL queries per knowledge graph, one for entities and one for properties. They compute one row per item with five columns each: IRI, label, score, synonyms, and additional information. The IRI is unique and besides it only the label is required. The rows are ordered by descending score, with a higher score corresponding to more popular or more commonly used items.  [^8] In the indices the score is used to resolve tie-breaks between equally ranked items. Synonyms define alternative labels that are indexed like the main label. [^9] When querying an index, labels and synonyms of each IRI are ranked separately, but the highest-ranked one determines the IRI’s final rank. The additional information is not indexed but can help to disambiguate similarly labeled items. See Table 2 for example data. [^10] See the corresponding SPARQL query at qlever.cs.uni-freiburg.de/wikidata/InnqDm.
@@ 4.1 Benchmarks and Metric
In the following, we introduce the main benchmark datasets we evaluate on in this work. We chose them because they are either widely used in the research community or very relevant to our work. We also evaluate our approach on more knowledge graphs and benchmarks, see Section 4.3 and Table 9 for details.
@@ 4.1.2 Metric
All models are evaluated using the F1-score averaged across samples. We adopt SPINACH’s modification, which permits additional columns (e.g., labels) in the predicted results without penalty. [^12] For results exceeding 1,024 rows, we revert to the standard exact F1-score due to the computational cost of row-wise assignment. Samples with empty groundtruth are excluded. If the predicted query is an ASK query and the groundtruth is a SELECT query (or vice versa), we assign a score of 1 if their results are semantically equivalent. To save compute and money, we evaluate on 200 random samples for each benchmark, which for QALD-7 and SPINACH is still the whole test set. For the component studies in Section 4.3, we use a subset of the benchmarks for the same reasons. [^13] Except for the function set study, were full coverage is key for a clear picture. Wherever we have access to the original predictions from previous work, we re-evaluate them with our metric and, for a fair comparison, only use those samples present in both our and their evaluation.
@@ 4.2 Results
Table 4 shows our zero-shot and few-shot results for GRASP and GRASP-F on the main Wikidata and Freebase benchmarks using GPT-4.1 models [32]. In the zero-shot setting we reach very good performance overall. The GPT-4.1 variant sets a new state-of-the-art on all Wikidata benchmarks, while coming close to or surpassing the best few-shot approaches on Freebase. The inclusion of feedback improves performance across the board: it thus helps if the model gets the chance to reassess and refine its own outputs once in a while. In the few-shot setting we only evaluate GPT-4.1 mini due to cost. As expected, adding examples improves performance even further, with few-shot GRASP using GPT-4.1 mini surpassing zero-shot GRASP-F using GPT-4.1 on 3 out of 5 benchmarks. Interestingly, feedback does not seem to help as much in the few-shot setting, probably because the in-context examples already provide sufficient information in this case.
@@ 4.3.1 Model Provider Study
Table 5 shows the results for GRASP using various base models. We find that open-source models like Qwen2.5 72B and Qwen3 32B [48] can reach performance equal to or better than commercial models like Gemini 2.0 Flash [19]. Most commercial models still outperform the open-source models, with GPT-4.1 and o4-mini [33] at the top. The reasoning models do not significantly outperform non-reasoning models, probably because we explicitly instruct GRASP to “think” before and after each step, which leads to reasoning-like behavior.
@@ 4.3.2 Function set study
We run Qwen2.5 72B on all benchmarks with each of our function sets. As expected, we find that the base function set $B$ without explicit search functions performs worst across all benchmarks. This is because the builtin search functionality in SPARQL on literals is limited compared to our search indices, often leads to errors due to high memory usage or timeouts, and is harder to use for the model. We find that overall the $S_{E}$ function set performs the best. We hypothesize that this is because it is easier to use than $S_{A}$ or $S_{C}$ while still allowing constrained search unlike $S$. See Table 6 for full results.
@@
CWQ WQSP QALD-10 QALD-7 SPINACH WWQ
@@ 4.3.3 Model Size Study
The Qwen2.5 and Qwen3 model families provide a range of different model sizes which we use to determine how model size affects performance with GRASP. As expected, we find that performance increases with model size. There also seems to be quite a jump in performance between the 7 B and 14 B models for Qwen2.5 and between the 4 B and 8 B models for Qwen3. The best model overall is Qwen2.5 72 B , though a model of comparable size is missing for Qwen3. See Table 8 for full results.
@@ 4.3.4 Few-shot and Feedback Study
In addition to the results from Table 4, we also run GPT-4.1 mini with random few-shot examples. As mentioned in Section 4.2, our results show that feedback generally improves performance if no examples are given but does not help as much, and sometimes even hurts, if there are examples. Also as expected, in most cases both types of examples improve performance, with similar examples helping more than random examples, especially on benchmarks with more similar train and test splits. We validate these findings with a separate set of evaluations using Qwen2.5 72B and observe similar patterns there. As a rule of thumb, one should generally use feedback for best performance, but it can be disabled if there are examples available, especially if they are known to be similar to expected future questions. See Table 8 for full results.
@@ 4.3.5 Knowledge Graph Study
To show that our approach also generalizes to other, less commonly evaluated knowledge graphs, we collect or create benchmarks for DBLP, DBpedia, ORKG, OSM, and UniProt, and add two more benchmarks for Wikidata and one for Freebase. We evaluate GRASP-F using GPT-4.1 on all of them and show the results in Table 9. Our approach performs well overall, except on OSM and UniProt (for reasons partly explained next).

### v2

# GRASP: Generic Reasoning And SPARQL Generation across Knowledge Graphs

@@ 1 Introduction
SELECT ?result WHERE {
?result p:P39 ?m .
?m ps:P39 wd:Q17989863 ; pq:P580 ?start .
OPTIONAL { ?m pq:P582 ?end . }
FILTER(?start <= "2011-12-31"8sd:date &&
(!BOUND(?end) ‘||‘ ?end >= "2011-12-31"8sd:date))
@@ 3.1 GRASP
With GRASP, the generation of SPARQL queries relies on a LLM that interacts with a knowledge graph autonomously by interleaving function calls to search or query the knowledge graph with reasoning about previous and future steps. Note, that this is in style similar to ReAct prompting [50], though we also allow multiple successive reasoning steps or function calls. Starting point of a generation is the GRASP instruction, which contains the name and endpoint of all available knowledge graphs, details about the step-by-step approach, and additional rules to be followed. Note, that this instruction is generic and does not contain any graph-specific information other than their names and endpoints. Optionally, few-shot examples (see Section˜3.5) and a feedback mechanism for refining the output (see Section˜3.6) can be added. See Fig.˜1 for an overview over GRASP’s approach by means of an example for DBLP.
@@ 3.3 Function Sets
[…]
@@ 3.4 Search Indices
The search functions are all based on pre-computed index data structures, which we describe in the following. See Table˜3 for details about all used indices.
@@ 3.4.1 Data
The data for the index data structures comes from two SPARQL queries per knowledge graph, one for entities and one for properties. They compute one row per item with five columns each: IRI, label, score, synonyms, and additional information. The IRI is unique and besides it only the label is required. The rows are ordered by descending score, with a higher score corresponding to more popular or more commonly used items.  [^8] In the indices the score is used to resolve tie-breaks between equally ranked items. Synonyms define alternative labels that are indexed like the main label. [^9] When querying an index, labels and synonyms of each IRI are ranked separately, but the highest-ranked one determines the IRI’s final rank. The additional information is not indexed but can help to disambiguate similarly labeled items. See Table˜2 for example data. [^10] See the corresponding SPARQL query at qlever.cs.uni-freiburg.de/wikidata/InnqDm.
@@ 4.1 Benchmarks and Metric
In the following, we introduce the main benchmark datasets we evaluate on in this work. We chose them because they are either widely used in the research community or very relevant to our work. We also evaluate our approach on more knowledge graphs and benchmarks, see Section˜4.3 and Table˜9 for details.
@@ 4.1.2 Metric
All models are evaluated using the F1-score averaged across samples. We adopt SPINACH’s modification, which permits additional columns (e.g., labels) in the predicted results without penalty. [^12] For results exceeding 1,024 rows, we revert to the standard exact F1-score due to the computational cost of row-wise assignment. Samples with empty groundtruth are excluded. If the predicted query is an ASK query and the groundtruth is a SELECT query (or vice versa), we assign a score of 1 if their results are semantically equivalent. To save compute and money, we evaluate on 200 random samples for each benchmark, which for QALD-7 and SPINACH is still the whole test set. For the component studies in Section˜4.3, we use a subset of the benchmarks for the same reasons. [^13] Except for the function set study, were full coverage is key for a clear picture. Wherever we have access to the original predictions from previous work, we re-evaluate them with our metric and, for a fair comparison, only use those samples present in both our and their evaluation.
@@ 4.2 Results
Table˜4 shows our zero-shot and few-shot results for GRASP and GRASP-F on the main Wikidata and Freebase benchmarks using GPT-4.1 models [32]. In the zero-shot setting we reach very good performance overall. The GPT-4.1 variant sets a new state-of-the-art on all Wikidata benchmarks, while coming close to or surpassing the best few-shot approaches on Freebase. The inclusion of feedback improves performance across the board: it thus helps if the model gets the chance to reassess and refine its own outputs once in a while. In the few-shot setting we only evaluate GPT-4.1 mini due to cost. As expected, adding examples improves performance even further, with few-shot GRASP using GPT-4.1 mini surpassing zero-shot GRASP-F using GPT-4.1 on 3 out of 5 benchmarks. Interestingly, feedback does not seem to help as much in the few-shot setting, probably because the in-context examples already provide sufficient information in this case.
@@ 4.3.1 Model Provider Study
Table˜5 shows the results for GRASP using various base models. We find that open-source models like Qwen2.5 72B and Qwen3 32B [49, 36] can reach performance equal to or better than commercial models like Gemini 2.0 Flash [19]. Most commercial models still outperform the open-source models, with GPT-4.1 and o4-mini [33] at the top. The reasoning models do not significantly outperform non-reasoning models, probably because we explicitly instruct GRASP to “think” before and after each step, which leads to reasoning-like behavior.
@@ 4.3.2 Function set study
[…]
@@
CWQ WQSP Q-10 Q-7 SPIN. WWQ
@@ 4.3.3 Model Size Study
The Qwen2.5 and Qwen3 model families provide a range of different model sizes which we use to determine how model size affects performance with GRASP. As expected, we find that performance increases with model size. There also seems to be quite a jump in performance between the 7 B and 14 B models for Qwen2.5 and between the 4 B and 8 B models for Qwen3. The best model overall is Qwen2.5 72 B , though a model of comparable size is missing for Qwen3. See Table˜8 for full results.
@@ 4.3.4 Few-shot and Feedback Study
In addition to the results from Table˜4, we also run GPT-4.1 mini with random few-shot examples. As mentioned in Section˜4.2, our results show that feedback generally improves performance if no examples are given but does not help as much, and sometimes even hurts, if there are examples. Also as expected, in most cases both types of examples improve performance, with similar examples helping more than random examples, especially on benchmarks with more similar train and test splits. We validate these findings with a separate set of evaluations using Qwen2.5 72B and observe similar patterns there. As a rule of thumb, one should generally use feedback for best performance, but it can be disabled if there are examples available, especially if they are known to be similar to expected future questions. See Table˜8 for full results.
@@ 4.3.5 Knowledge Graph Study
To show that our approach also generalizes to other, less commonly evaluated knowledge graphs, we collect or create benchmarks for DBLP, DBpedia, ORKG, OSM, and UniProt, and add two more benchmarks for Wikidata and one for Freebase. We evaluate GRASP-F using GPT-4.1 on all of them and show the results in Table˜9. Our approach performs well overall, except on OSM and UniProt (for reasons partly explained next).
@@ 5 Conclusion
5.0.1 Acknowledgments
Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 499552394 – SFB 1597.